Full price tickets at a concert is $56.50. Students can purchase discounted tickets for $ 49.50. The total tickets sold were 3150 and the total amount was $157,647. How many full priced tickets were sold? How many discounted tickets were sold?
Answer by mananth(16946)

Full price x numbers
student tickets y numbers

x+y=3150 .............1
56.5x+49.5y=157647 .............2
Eliminate y
multiply (1)by -49.5
Multiply (2) by 1
-49.5x-49.5y=-155925
56.5x+ 49.5y= 157647
Add the two equations
7x=1722
/ 7
x=246
plug value of x in (1)
x+y=3150
246+y=3150
y=3150-246
y=2904
Full price 246 numbers
student tickets 2904 numbers
m.ananth@hotmail.ca